how many numbers greater 250 can be formed using the digits 1,2,3,4 if no digit is to be repeated
Found 2 solutions by greenestamps, Alan3354:
Answer by greenestamps(13195)



(1) 4-digit numbers

All of the 4-digit numbers that can be made with those four digits are greater than 250. The number of them is 4*3*2*1 = 24.

(2) 3-digit numbers

The 3-digit numbers greater than 250 must have 3 or 4 as the first digit. The number of 3-digit numbers greater than 250 is 2*3*2 = 12.

Total number of numbers greater than 250 that can be formed using the digits 1, 2, 3, and 4 without repetition: 24+12 = 36.

Answer by Alan3354(69443)

how many numbers greater 250 can be formed using the digits 1,2,3,4 if no digit is to be repeated
---------
All the 4 digit #'s are > 250.
--> 4*3*2*1 = 24 4 digit #'s
-----
3 digit #'s must have 3 or 4 as the MSD.
--> 2*3*2*1 = 12 3 digit #'s
============
24 + 12 = 36